OpenStudy (anonymous):

solve for x: x^2/3 +5x^1/3 +6=0

OpenStudy (anonymous):

x^2/3 -5x^1/3 +6=0 %28x%5E%281%2F3%29%29%5E2-5x%5E%281%2F3%29%2B6=0 Let P=x^(1/3) P%5E2-5P%2B6=0 %28P-2%29%28P-3%29=0 p-2=0 or P-3=0 P-2+2=0+2 or P-3+3=0+3 P=2 or P=3 Let x^(1/3)=P x%5E%281%2F3%29=2 or x%5E%281%2F3%29=3 %28x%5E%281%2F3%29%29%5E3=2%5E3 or %28x%5E%281%2F3%29%29%5E3=3%5E3 highlight%28x=8%29 or highlight%28x=27%29 Check your answers: %288%29%5E%282%2F3%29-5%288%29%5E%281%2F3%29%2B6=0 %284%29-5%282%29%2B6=0 4-10%2B6=0 0=0 8's a right answer %2827%29%5E%282%2F3%29-5%2827%29%5E%281%2F3%29%2B6=0 9-5%283%29%2B6=0 9-15%2B6=0 0=0 27's right too!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

welcome!